# Automated Evaluation of Meter and Rhyme in Russian Generative and Human-Authored Poetry

### v1

@@ Abstract
In this work, we introduce the Russian Poetry Scansion Tool library designed for stress mark placement in Russian-language syllabo-tonic poetry, rhyme detection, and identification of defects of poeticness. Additionally, we release RIFMA – a dataset of poem fragments spanning various genres and forms, annotated with stress marks. This dataset can be used to evaluate the capability of modern large language models to accurately place stress marks in poetic texts.
@@ 1 Introduction
  • A new test dataset (7) comprising approximately 3,600 Russian human-authored poetry stanzas, each annotated with stress markings and rhyme scheme information. This dataset serves as a valuable resource for both research and benchmarking.
@@ 4 Russian Poetry Scansion Tool
There are two primary use cases for RPST. First, during data engineering, RPST can filter out defective poems with poorly observed meter and rhyme from the training dataset. This scenario is detailed in Section 6. Second, for evaluating generated poems, RPST provides a quantitative measure of technicality, enabling the ranking of outputs before presenting them to the user. This scenario is discussed in Section 5.
@@
5 Side-by-side Evaluation of Poetry

### v2

@@ Abstract
In this work, we introduce the Russian Poetry Scansion Tool library designed for stress mark placement in Russian-language syllabo-tonic poetry, rhyme detection, and identification of defects of poeticness. Additionally, we release RIFMA – a dataset of poem fragments spanning various genres and forms, annotated with stress marks and rhyme schemes. This dataset can be used to evaluate the capability of modern LLMs to accurately place stress marks in poetic texts.
@@ 1 Introduction
  • A new test dataset (7) comprising approximately 5,100 Russian human-authored poetry stanzas, each annotated with stress marks and rhyme scheme information. This dataset serves as a valuable resource for both research and benchmarking.
@@ 4 Russian Poetry Scansion Tool
There are two primary use cases for RPST. First, during data engineering, RPST can filter out defective poems with poorly observed meter and rhyme from the training dataset. This scenario is detailed in Section 6. Second, for evaluating generated poems, RPST provides a quantitative measure of technicality, enabling the ranking of outputs before presenting them to the user. This scenario is discussed in Sections 5.1 and 5.2.
@@
5 Human Evaluation
5.1 Human Assessment of Generative Poetry
We used the results of a crowdsourced assessment of rhyme and meter in Russian-language poems generated by several large language models (LLMs). In total, 875 generated texts were evaluated. Most texts were rated by five non-expert annotators, while a small portion received four or three ratings. For consistency, 859 texts with a complete overlap of five ratings were included in the subsequent analysis.
Each text was assigned a score on a three-point scale: 0 (no rhyme, meter not observed), 1 (partial rhyme or meter), and 2 (rhyme present, meter observed). The distribution of assigned scores was as follows:
  • Score 0: 952 ratings
  • Score 1: 1571 ratings
  • Score 2: 1772 ratings
For correlation analysis, the mean score of each text was computed across all raters. Each text was then analyzed using the RPST library, resulting in two corresponding lists of real-valued outputs. The correlation between these lists was computed using the scipy library.
The correlation is statistically significant at $\alpha=0.05$ (Pearson’s $r=0.79$, $p=3.2\times 10^{-181}$).
5.2 Side-by-side Evaluation of Poetry